OpenStudy (anonymous):

PLEASE HELP ASAP!!! A line has a y-intercept of 8 and a slope of 12/7. Find the equation of the line. y = 12/7x + 8 y = 8x - 7/12 y = 7/12x - 8 y = 8x + 12/7

OpenStudy (marissalovescats):

y=mx+b Where m is your slope And b is your y intercept

OpenStudy (marissalovescats):

So you just put 12/7 and 8 in the corresponding spots.

OpenStudy (anonymous):

its the first one.

OpenStudy (anonymous):

it was a could u help me with another? A line has a y-intercept of -2 and a slope of 8. Find the equation of the line. y = 8x - 16 y = 8x - 2 y = 8x + 2 y = -2x - 8

OpenStudy (marissalovescats):

It probably would of been better if he answered so he can plug things in and understand, but yes it is the first one. And this question i the same exact thing. Plus in the -2 and 8 where they go based on slope and y intercept.

OpenStudy (anonymous):

second one

OpenStudy (marissalovescats):

Telling him the answers is not beneficial to him. You want him to be able to do it on his own outside OpenStudy.